OpenStudy (kmullis6):

PLEASE HELP ME, ILL FAN & GIVE MEDAL (: Decide if lines p(1) and p(2) are perpendicular.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle y_1=\color{red}{\rm m} \cdot x+a}\) \(\large\color{black}{ \displaystyle y_2=\color{red}{\frac{-1}{\rm ~m}} \cdot x+b}\) then \(\large\color{black}{ \displaystyle y_1}\) and \(\large\color{black}{ \displaystyle y_2}\) are prependicular.

OpenStudy (kmullis6):